A father is three times as old as the son. Nine years ago the father was four times as old as the son find the present ages
Answer by ikleyn(52782)

.
A father is three times as old as the son. Nine years ago the father was four times as old as the son find the present ages.
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

F = 3S,          (1)   ("A father is three times as old as the son")
F-9 = 4*(S-9)    (2)   ("Nine years ago the father was four times as old as the son")

Substitute (1) into (2). You will get

3S-9 = 4(S-9),

3S-9 = 4S - 36   --->  S = 27.

The son is 27 years old now. The father is 3*27 = 81 years old now.

Answer.  The son is 27 years old now. The father is 81 years old now.

There is a bunch of lessons on age word problems
    - Age problems and their solutions
    - A fresh formulation of a traditional age problem
    - Really intricate age word problem
    - Selected age word problems from the archive
    - Age problem for the day of April, 1
    - OVERVIEW of lessons on age problems
in this site.

Read them and become an expert in solving age problems.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Age problems".